OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day?

OpenStudy (gorv):

2|x-6|=38-14=24 |x-6|=12 x-6=-12 or 12 x=-12+6 or 12+6 x=-6 or 18 min =-6 and max = 18

OpenStudy (anonymous):

2Ix-6I = 24 Ix-6I=12 now two equations... x-6 = -12 and x-6 =12 ...solve the two...

OpenStudy (gorv):

@harleeekoon

OpenStudy (nurali):

2|x − 6| + 14 = 38 2|x − 6| = 38-14 2|x − 6| =24 |x − 6|=24/2 |x − 6|=12 x-6=12 or x-6=-12 x=18 or x=-6

OpenStudy (anonymous):

but what were this the chose to this a. x = -9, x = 21 b. x = -6, x = 18 c. x = 6, x = -18 d. No solution